OpenStudy (anonymous):

CAN SOME ONE PLZ PLZ PLZ HELP ME WIT GOAL 4 ALGEBRA PLZ

OpenStudy (anonymous):

Sure (: Whats up?

OpenStudy (anonymous):

Thankk yu ok lyke for an example the problemz on my eoc practice sheet they give u an equation y=58.7(1.03)^t and t represents the number of years since january 1994 wat iz d propulatin of d country in 2002 wat kinda problem is that how du u solve them

OpenStudy (anonymous):

Okay so you have to subtract 1994 - 2002 = 8 so you plug eight in for t which give you 58.7 (1.03)^8 So then you have to.. hold on typing up more... ^_^ i did all this pretty recently ;D

OpenStudy (anonymous):

So then you have to do 1.03 *1.03*1.03*1.03*1.03*1.03*1.03*1.03 = 1.266.. So then you multiply by 58.7 1.266*58.7 = 74.3142 So thats the answer

OpenStudy (anonymous):

Does that help? (:

OpenStudy (anonymous):

Yesthankk yuu soe muchi really preciate ur help

OpenStudy (anonymous):

Welcome!! I am glad i helped (: